OpenStudy (anonymous):

Write the expression in a + bi form: (-√3 + i)^6

OpenStudy (ranga):

Have you been taught DeMoivre's Theorem or Binomial Theorem?

OpenStudy (ranga):

DeMoivre's Theorem: If z = r(cos(theta) + isin(theta), then z^n = r^n(cos(n*theta) + isin(n * theta)) Have you seen this formula before?

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (anonymous):

:)

OpenStudy (ranga):

If you have been taught DeMoivre's Theorem then I can tell you the steps to solve this.

OpenStudy (anonymous):

Okay i have

OpenStudy (anonymous):

@ranga why do you keep saying you will help me then just leave...

OpenStudy (ranga):

Because I posted my first reply 21 hours ago and you wouldn't even bother to answer the question that I had to reask the question again now.

OpenStudy (anonymous):

Oh im sorry i was waiting for a reply and then i went ofline

OpenStudy (ranga):

We need to put (-√3 + i) in the form r * ( cos(theta) + isin(theta) ) Find the magnitude of (-√3 + i). Factor it out. Magnitude of a + ib = sqrt(a^2 + b^2). I will wait for you to factor out the magnitude.

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Right?

OpenStudy (ranga):

Correct. Factor the magnitude out. (-√3 + i) = 2( -√3/2 + 1/2i ) Remember we need to put it in the form r * ( cos(theta) + isin(theta) ) Which angle theta will have cos(theta) = -√3/2 and sin(theta) = 1/2? (look up the unit circle).

OpenStudy (anonymous):

Okay hang on

OpenStudy (anonymous):

1/2 is 90 right?

OpenStudy (ranga):

sin(90) = 1, cos(90) = 0. Neither one gives half.

OpenStudy (anonymous):

oops sin(30)

OpenStudy (ranga):

sin(30) is 1/2 but cos(30) = +sqrt(3)/2. We want cos(theta) = -√3/2 and sin(theta) = 1/2. In a unit circle, the first coordinate is cos(theta) and the second coordinate is sin(theta). So look up unit circle and see which angle gives the coordinates (-sqrt(3)/2, 1/2).

OpenStudy (ranga):

The angle will be in the second quadrant because that is where cosine is negative and sine is positive.

OpenStudy (anonymous):

150!! :)

OpenStudy (ranga):

Yes! That is in degrees. Here it will be easier if we use radians. 150 degrees = 5*pi/6.

OpenStudy (ranga):

\[(-\sqrt{3} + i) = 2(-\sqrt{3}/2 + 1/2i) = 2[ \cos(5*\pi/6) + isin(5*\pi/6) ]\]

OpenStudy (ranga):

\[(-\sqrt{3} + i)^6 = [2( \cos(5*\pi/6) + isin(5*\pi/6) )]^6 = \]

OpenStudy (ranga):

\[2^6 * ( \cos(5*\pi) + isin(5*\pi) ) = 64 * (\cos(\pi) + isin(\pi)) = 64 * -1 = -64\]

OpenStudy (anonymous):

Okay still trying to understand all of it but why did the 5 just disappear?

OpenStudy (ranga):

In a unit circle, 2pi is one full rotation around a circle and so any additions or subtraction of 2*pi from the angle will land up in the same place. 5*pi = pi + 2pi + 2pi. We can subtract the two 2pi's because they are just two full rounds around the unit circle and you end up at the same angle.

OpenStudy (anonymous):

Ohhh that makes sense :)

OpenStudy (ranga):

cos(5pi) is same as cos(3pi) is same as cos(pi) is same as cos(7pi), cos(9pi), etc.

OpenStudy (anonymous):

Hmm never knew that but i guess it makes sense because of what you said about 2pi :)

OpenStudy (anonymous):

Can you help me with another similar one except it says to put it in a + bi form? ill open up a new question if you can help?

OpenStudy (ranga):

In degrees you can keep adding multiples of 360 degrees or in radians multiples of 2pi and ALL trig function will be unaltered. sin(30) = sin(390) = sin(750) (where I keep adding 360 degrees). tan(10) = tan(370) = tan(730), etc.. okay. post it as a separate question.

OpenStudy (anonymous):

OKay thanks thats helpful

OpenStudy (ranga):

You are welcome. Close this question and post a new one.